A parallel-plate capacitor with plate area 20 cm^2 and plate separation 1.0 mm is connected to a battery. The resistance of the circuit is 10 k Ω. Find the time constant of the circuit.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. A = 20 cm2 = 20 * 10^–4 m2 d = 1 mm = 1 * 10–3 m ; R = 10 KΩ C = E base 0A/d = 8.85 * 10^-12 * 20 * 10^-4/1 * 10^-3 = 8.85 * 10^-12 * 2 * 10^-3/10^-3 = 17.7 * 10^-2 Farad. Time constant = CR = 17.7 * 10^–2 * 10 * 10^3 = 17.7 * 10^–8 = 0.177 * 10^–6 s = 0.18 μs.